Amina has saved $350,000 for retirement, and her account earns 5% interest. How much can she pull out of her account each month, if she wants to be able to take withdrawals for 20 years?
Answer #1

PV = 350,000
Rate = 5%/12
Number of Periods = 20*12 = 240
PMT = 350,000/(1-(1+5%/12)-240)/5%/12 = 2309.85

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well